x-one whole number three over five 1 3/5 = two whole number seven over ten 2 7/10
Answer by checkley75(3666)

x-(1 3/5)=(2 7/10)
x-8/5=27/10
x=27/10+8/5 now we need to find a common denominator so we can combine these fractions thus: Here the common denominator is 10.
x=27/10+16/10 now we can add these fractions thus:
x=(27+16)/10
x=43/10
x=4.3